2 liters of white wine that is 12% alcohol. How many liters of white grape juice should be added to get a drink that is 10% alcohol.
Answer by stanbon(75887)

2 liters of white wine that is 12% alcohol. How many liters of white grape juice should be added to get a drink that is 10% alcohol.
----------------------
Equation:
alch + alch = alch
0.12*2 + 0*x = 0.10(2+x)
-------------------------------
12*2 + 0 = 10*2 + 10x
----
10x = 2*2
x = (2/5) = 0.4 liters (amt. of white grape juice to add)
---------------------
Cheers,
Stan H.
--------------